OpenStudy (anonymous):

If g(x) = 3x + 1, evaluate g(-14).

OpenStudy (bibby):

Plug the value into the equation and evaluate. what is \[(3*-14) + 1\]

OpenStudy (anonymous):

waaa??? lol

OpenStudy (bibby):

what is \[3 \times -14\]

OpenStudy (anonymous):

11?

OpenStudy (bibby):

that's a multiplication sign. Can you not multiply?

OpenStudy (anonymous):

-42 my bad lol

OpenStudy (bibby):

lol I sounded annoyed. I am not annoyed dw. ok so we have \[g(-14) = (3*-14) + 1\]\[g(-14) = -42 + 1\]

OpenStudy (anonymous):

so we solve now or what? lol i just want the answr my friend not tryna be speedy about it but i gotta get 10 credit recoverys done and i only got one done): and it alg 2):

OpenStudy (bibby):

solve the damn equation dude Can you not add either?

OpenStudy (anonymous):

mean retrice!!!!!!! yeah i can but im not on top my game right now

OpenStudy (anonymous):

retrice*

OpenStudy (bibby):

Pretend I'm your coach that's extra tough but just loves you but not in that way

OpenStudy (bibby):

add -42 and 1. I'm not sure what you're trying to do

OpenStudy (anonymous):

thanks my dude! but now i got alot more will you help me plz!

OpenStudy (bibby):

uh yeah. Just close your question, give medals, post new questions. Repeat until we clear your assignment